# VLA-Cache: Towards Efficient Vision-Language-Action Model via Adaptive Token Caching in Robotic Manipulation

### v1

@@ 4.1 Dynamic Token Selection
where $v_{\text{start}},v_{\text{end}}$ and $t_{\text{start}},t_{\text{end}}$ are the indices of the vision and text tokens, respectively.
To aggregate the attention scores across multiple heads, we compute the mean attention for each vision token as $\mathbf{A}^{l}_{\text{avg}}=\mathrm{Mean}_{\text{heads}}\!\bigl{(}\mathbf{A}^{l}_{\text{vis-text}}\bigr{)}$. For task relevance across multiple layers $\mathcal{L}$, the final task relevance scores are obtained by averaging the scores across the selected layers as $\mathbf{S}_{\text{task-relevance}}=\mathrm{Mean}_{l\in\mathcal{L}}\!\bigl{(}\mathbf{A}^{l}_{\text{avg}}\bigr{)}$.
Using these scores, we rank the vision tokens based on their task relevance and apply a threshold $\tau_{\text{task}}$ to select the most task-relevant tokens:
@@
$\mathcal{P}_{\mathrm{final}}=\mathcal{P}_{\mathrm{static}}\;\setminus\;\mathcal{P}_{\text{task-relevant}}.$ (10)
By filtering out high-attention tokens, this approach ensures that the VLA model accurately updates the critical objects in the scene, thereby preventing failures caused by reusing outdated static tokens. This method effectively balances computational efficiency with task precision, enabling robust performance in robotic control tasks.
4.2 Layer Adaptive Token Reusing
While static token selection and task-relevance filtering eliminate a large portion of redundant computation, we observe that attention distributions within the VLA decoder vary significantly across different layers. This finding is consistent with observations reported by prior work FastV (Chen et al., 2025), indicating that both VLA and VLM decoders exhibit similar patterns of attention flow: early layers display dispersed attention, followed by fluctuations in intermediate layers, and eventually a partial rebound near the final layers.
To account for these differences, we propose a layer-adaptive strategy that adjusts the fraction of reused tokens based on each layer’s attention concentration. Specifically, we quantify the attention distribution at layer $l$ via an entropy measure, following the same mean-attention computation described in equation  8. Let $H^{l}$ denote the resulting entropy. We then define an entropy ratio $R^{l}$, which captures how much more concentrated the attention is in layer $l$ compared to layer $l-1$:
$R^{l}\;=\;\frac{H^{l-1}-H^{l}}{H^{l-1}}.$ (11)
A positive $R^{l}$ indicates that the attention distribution at layer $l$ is more focused than that of layer $l-1$.
We accumulate these ratios across layers to obtain a cumulative score $R_{\text{cum}}^{l}=\sum_{j=1}^{l}R^{j}$, which in turn determines the proportion $\alpha^{l}$ of static tokens (from $\mathcal{P}_{\mathrm{final}}$) that are reused at layer $l$. Formally,
@@ 4.2 Layer Adaptive Token Reusing
Algorithm 1 Adaptive Token Caching
1:  Input: Frames $\{I_{t-1},I_{t}\}$, previous KV cache $\{\mathbf{K}_{t-1}^{l},\mathbf{V}_{t-1}^{l}\}$, thresholds $\tau,\tau_{\text{task}}$, hyperparameter $k$
2:  Output: Updated KV cache $\{\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}\}$
3:  Static Token Selection:
4:  Patchify $I_{t-1},I_{t}$ and compute $\text{Sim}(\mathbf{P}_{t},\mathbf{P}_{t-1})$
@@
11:  Layer-Adaptive Token Reuse:
@@
12:  Compute entropy reduction $R^{l}$, $R_{\text{cum}}^{l}$; Determine reuse ratio $\alpha^{l}$; Compute reusable subset $\mathcal{P}_{\mathrm{reuse}}\subseteq\mathcal{P}_{\mathrm{final}}$
13:  for each layer $l$ and token $i$ do
14:     if $i\in\mathcal{P}_{\mathrm{reuse}}$ then
15:        Reuse cached values:    $\mathbf{K}_{t}^{l}(i)=\mathbf{K}_{t-1}^{l}(I),\mathbf{V}_{t}^{l}(i)=\mathbf{V}_{t-1}^{l}(i)$
16:     else
17:        Recompute:    $\mathbf{K}_{t}^{l}(i)=W_{K}^{l}\mathbf{H}_{t}^{l}(i),\mathbf{V}_{t}^{l}(i)=W_{V}^{l}\mathbf{H}_{t}^{l}(i)$
18:     end if
19:     Compute self-attention:
20:     $\mathrm{Attn}(\mathbf{Q}_{t}^{l},\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l})=\mathrm{Softmax}\!\Bigl{(}\frac{\mathbf{Q}_{t}^{l}(\mathbf{K}_{t}^{l})^{\top}}{\sqrt{D}}\Bigr{)}\mathbf{V}_{t}^{l}$
21:  end for
22:  return Updated KV cache $\{\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}\}$
Inference with Cached Representation. During inference, VLA-Cache selectively caches and reuses previously computed representations to reduce redundant computations while maintaining accurate robotic action prediction. Instead of recomputing all tokens at each timestep, the model distinguishes between static and dynamic tokens. Static tokens, which exhibit minimal visual change across consecutive frames, retain their cached key-value (KV) representations from the previous timestep, bypassing unnecessary computation. Dynamic tokens, which undergo significant changes or are task-relevant, are recomputed to ensure precise motion execution. This selective token reuse is performed at each layer, guided by entropy-based attention analysis to dynamically adjust reuse ratios. The full inference procedure, including static token selection, task-relevance filtering, and layer-adaptive token reuse, is outlined in Algorithm 1, with further details provided in Appendix A.2.
@@ 4.3 Theoretical Analysis of Computational Complexity
Our method reduces the effective token count per layer to $L_{r}=\alpha\times\mathcal{P}_{\mathrm{final}}$, leading to theoretical savings:
@@
These savings scale across multiple layers, significantly lowering computation while maintaining accuracy.
@@
$\Delta\text{FLOPs}_{\text{total}}\approx$ $\Bigl{(}4L_{r}D^{2}+2L_{r}^{2}D+2L_{r}DM\Bigr{)}$ (15)
$-\Bigl{(}H^{2}+L_{t}L_{v}D+L^{2}D\Bigr{)}.$
Further Derivations and Complexity Details. See Appendix A.1 for detailed derivations, including static token selection costs, attention filtering complexity, and layer-adaptive entropy calculations.
@@ 5 Experiment
5.1 Experimental Settings
To validate the generalization of VLA-Cache, we implement our method in both simulation and real-world settings. In simulation, we evaluate VLA-Cache on two open-source VLA models: OpenVLA (Kim et al., 2024) and CogAct (Li et al., 2024a), using the LIBERO benchmark (Liu et al., 2024a) and SIMPLER environment (Li et al., 2024b), respectively. For fair comparison, we adopt two token-pruning acceleration methods, SparseVLM (Zhang et al., 2024) and FastV (Chen et al., 2025), as baselines in the LIBERO environment. OpenVLA processes 256 visual tokens per image, and all comparative experiments use an identical token pruning and reuse count, setting Top-K to 100 tokens.
For VLA-Cache, the default parameters are set as follows: similarity score threshold $\tau=0.996$, top-$k=100$, and task relevance threshold $\tau_{task}=0.5$. Additionally, we use OpenVLA’s default caching mechanism and other default parameters for inference acceleration. In the SIMPLER environment, we apply VLA-Cache with the same parameter settings as OpenVLA while using CogAct’s default hyperparameters.
For real-world evaluation, we fine-tune OpenVLA for deployment on a Kinova Jaco2 6-DoF manipulator. The model is trained for 50,000 steps with default parameters. During testing, we modify only the similarity score threshold to $\tau=0.85$, keeping all other parameters unchanged. All experiments are conducted on an NVIDIA RTX 4090 GPU.
5.1.1 LIBERO Benchmark
The LIBERO Benchmark (Liu et al., 2024a) is designed to evaluate lifelong robotic manipulation across four diverse task suites: LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long. Each suite introduces distinct variations in spatial layouts, object selection, task objectives, or long-horizon planning, challenging models to generalize across different manipulation scenarios.
Following the experimental setup of OpenVLA, we conduct evaluations on all four LIBERO suites, each containing ten subtasks. To ensure consistency, we use the same machine and official fine-tuned weights. For each subtask, multiple evaluation episodes are run, measuring key performance metrics such as success rate and inference latency. We evaluate VLA-Cache’s acceleration capabilities within the OpenVLA framework and conduct an ablation study on the top-$k$ parameter in the LIBERO environment.
Figure 3: Simulation Tasks on LIBERO Benchmark and the SIMPLER Environment.
Figure 4: Robot Manipulation Tasks on Jaco2 Real-world Platform.
Figure 5: Heatmap Visualization of Robot Actions in a Simulated Environment with Prompt Descriptions and Outputs
[…]
LIBERO-Spatial LIBERO-Object LIBERO-Goal LIBERO-Long Average
Baseline (OpenVLA) 84.4% 86.6% 75.6% 53.2% 75.0% 1.864 51.91
SparseVLM 79.8% 67.0% 72.6% 39.4% 64.7% 1.407 83.39
FastV 83.4% 84.0% 74.2% 51.6% 73.3% 1.864 53.28
VLA-Cache 83.8% 85.8% 76.4% 52.8% 74.7% 1.355 31.83
@@ 5.1.1 LIBERO Benchmark
Visual Matching Baseline (CogACT) 91.3% 85.0% 71.8% 50.9% 74.8% 1.847 54.29
VLA-Cache 92.0% 83.3% 70.5% 51.6% 74.4% 1.496 39.63
Variant Aggregation Baseline (CogACT) 89.6% 80.8% 28.3% 46.6% 61.3% 1.807 53.54
VLA-Cache 91.7% 79.3% 32.5% 45.8% 62.3% 1.493 39.11
Table 2: Comparison of VLA Cache in CogACT model within the SIMPLER environment.
5.1.2 SIMPLER Environments
The SIMPLER simulation environment (Li et al., 2024b) is designed to closely mirror real-world robotic setups, facilitating more realistic evaluations.
SIMPLER supports two evaluation settings: Visual Matching (aiming to minimize visual discrepancies between simulation and reality) and Variant Aggregation (building on Visual Matching by introducing variations in lighting, background, table textures, etc.). We adopt the same SIMPLER configuration used by CogAct and evaluate two settings, Visual Matching and Variant Aggregation, on a Google robot arm. Each setting contains four tasks: 1) Pick coke can(PickCan), 2) Move near(MoveNear), 3) Open/close drawer(Drawer), 4) Open top drawer and place apple(DrawerApple).
To validate the general applicability of our method, we use CogAct as the baseline model within the SIMPLER environment. CogAct extends OpenVLA by incorporating a Diffusion Policy head to predict actions in continuous space, a common strategy among recent VLA models. Similar to OpenVLA, CogAct provides publicly available weights and code for reproducible simulation-based evaluations. Through these experiments, we try to verify its potential as a general VLA acceleration technique.
5.1.3 Real-world Robot Setting
We further evaluate our method on a Kinova Jaco2 6-DoF manipulator, equipped with a Sony AX53 camera positioned in front of the arm to capture the manipulation scene.
Tasks. We design four real-world robotic manipulation tasks:1) Pick up the orange pot (PickPot),2) Place the blue cube in the box (PlaceCube), 3) Put the sausage in the blue pan (PutSausage),4) Wipe the table (WipeTable). For the wiping task, we place objects of various shapes and colors on the table to challenge the model’s robustness to unseen objects.
Data Collection. Using an Xbox Gamepad, we teleoperate the Jaco2 arm and collect demonstrations at a frequency of 10 Hz, recording RGB images and robot states throughout each trajectory. We gather 150–200 demonstrations per task to form the training dataset.
Baseline. We fine-tune OpenVLA via LoRA (Low-Rank Adaptation (Hu et al., 2021)) on each task and then deploy it on the Jaco2 to assess real-world performance. The fine-tuning process follows the same hyperparameter setup as in LIBERO.
5.2 Results on Simulation Environment
@@
PickPot PlaceCube PutSausage WipeTable Average
Baseline (OpenVLA) 95.0% 83.3% 80.0% 70.0% 82.1% 1.814 64.16
VLA-Cache 90.0% 90.0% 85.0% 73.0% 84.5% 1.303 51.85
Table 3: Comparison of success rate on real robot tasks.
[…]
Static Token 74.2% 1.304 31.03
Task Relevant 82.6% 1.304 31.03
Layer Adaptive (Ours) 83.8% 1.382 32.22
Table 4: Comparison of different token selection methods.
  #Tokens Methods SR % $\uparrow$ FLOPs $\downarrow$ Time(ms) $\downarrow$
@@
Table 5: LIBERO-Spatial benchmark comparing different methods based on pruning or reuse tokens. Best values in each column are highlighted in bold.
Main Results on LIBERO.
Table 1 summarizes the success rate, FLOPs, and CUDA inference time across the four LIBERO task suites. Compared to the baseline, VLA-Cache achieves a 27.31% FLOPs reduction and a 1.63$\times$ speedup, with only a minimal 0.3% drop in the overall success rate, confirming its efficiency without compromising accuracy. While VLA-Cache slightly underperforms on spatial, object, and long-horizon tasks, it notably surpasses the baseline on the goal suite, highlighting its robust adaptability.
Notably, both FastV and SparseVLM fail to improve inference time, often exceeding the baseline. FastV merely masks tokens in attention computation without reducing GPU workload, introducing additional masking overhead, while SparseVLM incurs extra cost from token pruning, merging, and recycling. These methods were originally designed for long-sequence VLM tasks, where efficiency gains are more pronounced, but struggle in VLA settings, where action sequences are typically short. In contrast, VLA-Cache leverages inter-frame token reuse with negligible overhead, making it effective for robotic action generation.
Ablation on Reusing/Pruning Rate. Table 5 presents a comparative analysis of different methods through an ablation study on the top-$k$ parameter, while also varying the number of pruned or reused tokens. For all methods, increasing the number of reused/pruned tokens leads to a decline in success rates, highlighting the importance of retaining sufficient information for accurate action predictions. Our approach maintains a relatively stable success rate at a reuse rate of 40%, but a more noticeable drop occurs when reusing 200 tokens. In contrast, FastV removes a significant portion of tokens entirely, and its performance drop primarily results from the loss of critical visual details rather than potential errors in token reuse. Additionally, SparseVLM and FastV incur longer inference times due to extra attention operations during GPU execution, whereas VLA-Cache directly updates KV cache entries, making inference more efficient.
Token Selection Strategies As presented in Table 4, we evaluate different token selection strategies for VLA-Cache. The direct reuse of static tokens results in a significant decline in success rate. In contrast, filtering task-relevant tokens maintains a success rate of 82.6%. Notably, our layer-adaptive approach further enhances the success rate to 83.8%, albeit with a marginal increase in computational overhead from 1.304 FLOPs to 1.382 FLOPs. This overhead arises because fewer tokens are reused in the early layers, which ensures the retention of critical updates in deeper layers, thereby improving overall performance.
@@
Table 2 indicates that the VLA-Cache exhibits comparable success rates than the baseline CogACT in the SIMPLER environment while substantially reducing computational overhead. Specifically, VLA-Cache achieves an average success rate of 74.4% in the standard Visual Matching setting compared to 74.8% for CogACT and 62.3% in the Variant Aggregation setting versus 61.3% for CogACT.
The efficiency gains are evident in the FLOPs and inference time measurements. VLA-Cache achieves roughly 20% fewer FLOPs than the baseline, coupled with a 1.37× reduction in inference latency. Notably, these results highlight the portability of VLA-Cache across different action heads, establishing it as a general acceleration strategy for VLA.

### v2

@@ 3.3 Retaining Task-Relevant Information
where $v_{\text{start}},v_{\text{end}}$ and $t_{\text{start}},t_{\text{end}}$ are the indices of the vision and text tokens, respectively. To aggregate the attention scores across multiple heads, we compute the mean attention for each vision token as $\mathbf{A}^{l}_{\text{avg}}=\mathrm{Mean}_{\text{heads}}\!\bigl(\mathbf{A}^{l}_{\text{vis-text}}\bigr)$. For task relevance across multiple layers $\mathcal{L}$, the final task relevance scores are obtained by averaging the scores across the selected layers as $\mathbf{S}_{\text{task-relevance}}=\mathrm{Mean}_{l\in\mathcal{L}}\!\bigl(\mathbf{A}^{l}_{\text{avg}}\bigr)$. Using these scores, we rank the vision tokens based on their task relevance and apply a threshold $\tau_{\text{task}}$ to select the most task-relevant tokens:
@@
$\mathcal{P}_{\mathrm{reuse}}=\mathcal{P}_{\mathrm{static}}\;\setminus\;\mathcal{P}_{\text{task-relevant}}.$ (8)
By filtering out semantically significant tokens from the static reuse set, our method restores the degraded success rate from 74.2% to 82.6%, while maintaining the computational gains in FLOPs and latency. This balance between task fidelity and efficiency illustrates the value of cross-modal attention as a lightweight signal for safe token reuse in VLA models.
Algorithm 1 Dynamic Token Selection
1:Input: Frames $\{I_{t-1},I_{t}\}$, thresholds $\tau$, $\tau_{\text{task}}$, hyperparameter $k$
2:Output: Reusable token indices $\mathcal{P}_{\mathrm{final}}$
@@
4: Patchify both frames $I_{t-1},I_{t}$ and compute $\text{Sim}(\mathbf{P}_{t},\mathbf{P}_{t-1})$ among related patches
@@
11:return $\mathcal{P}_{\mathrm{final}}$
Algorithm 2 Adaptive Token Caching
1:Input: Token indices $\mathcal{P}_{\mathrm{final}}$, previous KV cache $\{\mathbf{K}_{t-1}^{l},\mathbf{V}_{t-1}^{l}\}$, representations $\mathbf{H}_{t}^{l}$
2:Output: Updated KV cache $\{\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}\}$
@@
3: Compute entropy $R_{\text{cum}}^{l}$, reuse ratio $\alpha^{l}$, subset $\mathcal{P}_{\mathrm{reuse}}\subseteq\mathcal{P}_{\mathrm{final}}$
4:for each layer $l$ and token $i$ do
5:  if $i\in\mathcal{P}_{\mathrm{reuse}}$ then
6:   Reuse cached values: $\mathbf{K}_{t}^{l}(i)=\mathbf{K}_{t-1}^{l}(i),\mathbf{V}_{t}^{l}(i)=\mathbf{V}_{t-1}^{l}(i)$
7:  else
[…]
9:  end if
10:end for
11:return $\{\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}\}$
3.4 Layer Adaptive Token Reusing
While static token selection and task-relevance filtering eliminate a large portion of redundant computation, we observe that attention distributions within the VLA decoder vary significantly across different layers. This finding is consistent with observations reported by prior work FastV [29], indicating that both VLA and VLM decoders exhibit similar patterns of attention flow: early layers display dispersed attention, followed by fluctuations in intermediate layers, and eventually a partial rebound near the final layers.
To account for these differences, we propose a layer-adaptive strategy that adjusts the fraction of reused tokens based on each layer’s attention concentration. Specifically, we quantify the attention distribution at layer $l$ via an entropy measure, following the same mean-attention computation described in Eq.  6. Let $\mathcal{E}^{l}$ denote the resulting entropy. We then define an entropy ratio $R^{l}=(\mathcal{E}^{l-1}-\mathcal{E}^{l})/\mathcal{E}^{l-1}$, which captures how much more concentrated the attention is in layer $l$ compared to layer l1l-1.
A positive $R^{l}$ indicates that the attention distribution at layer $l$ is more focused than that of layer l1l-1. We accumulate these ratios across layers to obtain a cumulative score, which in turn determines the proportion $\alpha^{l}$ of static tokens (from $\mathcal{P}_{\mathrm{final}}$) that are reused at layer $l$. Formally,
$\alpha^{l}\;=\;\min\Bigl(k\sum_{j=1}^{l}R^{j},\;1\Bigr),$ (9)
4 Implementations
4.1 Cross-Frame Visual Token Caching
During inference, VLA-Cache accelerates robotic action prediction by reusing previously computed key-value (KV) representations of visual tokens across time. Instead of recomputing all visual tokens in each time step, the model identifies a subset of static tokens, those that exhibit minimal change across frames, and reuses their cached representations from the previous time step. In contrast, dynamic tokens, which undergo significant visual change or are task-relevant, are freshly computed to maintain accurate action generation.
At each timestep $t$, given the visual token sequence $\mathbf{H}_{t}$, the model identifies a subset $\mathcal{P}_{\text{reuse}}$ of tokens that remain unchanged from the previous frame and reuses their cached representations:
$\mathbf{K}_{t}(i)=\begin{cases}\mathbf{K}_{t-1}(i),&i\in\mathcal{P}_{\text{reuse}}\\ W_{K}\mathbf{H}_{t}(i),&\text{otherwise}\end{cases},\quad\mathbf{V}_{t}(i)=\begin{cases}\mathbf{V}_{t-1}(i),&i\in\mathcal{P}_{\text{reuse}}\\ W_{V}\mathbf{H}_{t}(i),&\text{otherwise}\end{cases}.$ (10)
This design avoids redundant computation for static tokens while ensuring dynamic or task-relevant inputs are freshly computed. Most existing approaches reduce computation by pruning or merging tokens within a single frame [29, 30, 31, 32, 33]. In contrast, VLA-Cache exploits temporal redundancy by caching and reusing visual tokens across frames, making it better aligned with the closed-loop nature of robotic control. Notably, it is compatible with high-frequency architectures and directly alleviates the decoding bottleneck. Since it requires no model modification or retraining, VLA-Cache serves as a plug-and-play optimization for efficient robotic inference. An overview of the inference procedure is shown in Algorithm 1 and Algorithm 2, detailing token selection and adaptive caching. Additional implementation details are provided in Appendix D.
@@ 4.2 Theoretical Analysis of Computational Complexity
VLA-Cache reduces effective token count per layer to $L_{r}=\alpha\times\mathcal{P}_{\mathrm{final}}$, leading to theoretical savings:
@@
$\Delta\text{FLOPs}_{\text{total}}\approx\Bigl(4L_{r}D^{2}+2L_{r}^{2}D+2L_{r}DM\Bigr)-\Bigl(H^{2}+L_{t}L_{v}D+L^{2}D\Bigr).$ (13)
Please refer to Appendix C for detailed derivations, including static token selection costs, attention filtering complexity, and layer-adaptive entropy calculations.
@@ 5 Experiment
To validate the effectiveness of VLA-Cache, we evaluate our method in both simulation and real-world settings. In simulation, we evaluate VLA-Cache on three open-source VLA models: OpenVLA [11], OpenVLA-OFT [20] and CogAct [19], using the LIBERO benchmark [17] and SIMPLER environment [18], respectively. All experiments are conducted on an NVIDIA RTX 4090 GPU.
5.1 Experiment Setup
Compared Methods. We leverage the architectural similarity between VLA and VLM models, which allows direct application of existing VLM acceleration methods to VLA inference. Specifically, we adopt two state-of-the-art token-level acceleration techniques SparseVLM [30] and FastV [29] on OpenVLA as compared methods in the LIBERO benchmark.
Evaluation Metrics. We evaluate VLA-Cache using four metrics: success rate, control frequency, FLOPs, and CUDA latency. Success rate and control frequency respectively assess task performance and the responsiveness of action prediction in closed-loop control. FLOPs measure theoretical computation, while CUDA latency captures actual GPU runtime. These two efficiency metrics are widely adopted in VLM/VLA acceleration methods.
5.2 Evaluation Benchmark
LIBERO. The LIBERO Benchmark [17] covers four task suites: Spatial, Object, Goal, and Long, each testing a different aspect of manipulation generalization. We follow the standard setup from OpenVLA [11] and OpenVLA-OFT [20], using official weights and machines for consistency. Each suite includes ten subtasks evaluated over multiple episodes.
Figure 3: Tasks on LIBERO Benchmark, the SIMPLER Environment and Real World.
SIMPLER. The SIMPLER simulator [18] offers two settings, Visual Matching and Variant Aggregation, designed to bridge simulation-to-reality gaps. Following CogAct’s setup [19], we evaluate both settings on a Google robot arm across four manipulation tasks. CogAct, which integrates a Diffusion Policy for continuous control, serves as our baseline. These evaluations demonstrate the generality of VLA-Cache under different action heads and simulation variations.
Real Robot Evaluation. We deploy VLA-Cache on a Kinova Jaco2 manipulator equipped with a front-facing camera. The robot is evaluated on four tasks: PickPot, PlaceCube, PutSausage, and WipeTable, with the last task including diverse distractor objects to test robustness. Demonstrations are collected via teleoperation at 10 Hz using an Xbox controller, resulting in 150-200 trajectories per task. We fine-tune OpenVLA with LoRA [38] and evaluate on the same tasks, using the LIBERO tuning setup for consistency. More details about real-world experiments are available in the AppendixE.4.
5.3 Results on Simulation Environment
@@
Spatial Object Goal Long Average
OpenVLA 84.4% 86.6% 75.6% 53.2% 75.0% 1.864 51.91 4.23
+ SparseVLM 79.8% 67.0% 72.6% 39.4% 64.7% 1.407 83.39 3.72
+ FastV 83.4% 84.0% 74.2% 51.6% 73.3% 1.864 53.28 4.19
+ VLA-Cache 83.8% 85.8% 76.4% 52.8% 74.7% 1.355 31.83 4.59
OpenVLA-OFT 97.8% 97.6% 97.6% 94.2% 96.8% 4.013 79.05 65.10
+ VLA-Cache 98.3% 97.5% 98.3% 95.4% 97.4% 3.097 62.59 78.98
Table 3: Comparison of VLA-Cache within the CogACT model in the SIMPLER environment.
[…]
@@
Matching CogACT 91.3% 85.0% 71.8% 50.9% 74.8% 1.847 54.29 12.42
     + VLA-Cache 92.0% 83.3% 70.5% 51.6% 74.4% 1.496 39.63 14.66
Aggregation CogACT 89.6% 80.8% 28.3% 46.6% 61.3% 1.807 53.54 12.36
     + VLA-Cache 91.7% 79.3% 32.5% 45.8% 62.3% 1.493 39.11 14.48
Main Results on LIBERO.
@@
Table 4: Ablation on token pruning/reuse in LIBERO-Spatial using OpenVLA (256 vision tokens). Best values are in bold.
#Tokens Methods SR % $\uparrow$ FLOPs $\downarrow$ Latency (ms) $\downarrow$
@@
As illustrated in Figure 4, VLA-Cache effectively reduces visual computation redundancy during robotic manipulation by precisely identifying static tokens and filtering out task-relevant regions in real time. Notably, OpenVLA-OFT takes both fixed third-person and dynamic wrist-camera views as inputs. Its strong performance demonstrates that VLA-Cache not only enhances control frequency but also maintains robustness under dynamic viewpoints and camera shifts.
Ablation on Token Reusing/Pruning Rate. Table 4 presents an ablation study varying the number of reused/pruned tokens. For all methods, aggressive token reduction harms success rate, underscoring the need to preserve informative content. VLA-Cache maintains stable performance at moderate reuse rates (i.e., 100 tokens), while FastV and SparseVLM suffer larger drops due to loss of critical visual details. By directly updating KV entries, VLA-Cache remains more efficient and robust under different token reuse configurations.
Token Selection Strategies. As shown in Table 1, directly reusing all static tokens leads to a notable drop in success rate 74.2%, indicating that visual similarity alone is insufficient for reliable reuse in robotic control. By filtering out task-relevant tokens based on decoder attention, our method recovers performance to 82.6%. Introducing a layer-adaptive strategy further improves accuracy to 83.8%, with minimal increase in CUDA latency.
@@
Table 3 indicates that VLA-Cache exhibits success rates comparable to the CogACT baseline in the SIMPLER environment while substantially reducing computational overhead.
[…]
Figure 4: Visualization of VLA-Cache token reuse across settings. (a) LIBERO simulation with OpenVLA. (b) Real-world task under dynamic background. (c) and (d) Main and wrist camera views from OpenVLA-OFT. Blue: static tokens, Yellow: task-relevant, Red: overlapping. VLA-Cache reduces redundant computation and preserves accuracy under varying conditions.
Table 5: Comparison of success rate on real robot tasks.
[…]
PickPot PlaceCube PutSausage WipeTable Average
OpenVLA 95.0% 83.3% 80.0% 70.0% 82.1% 1.814 64.16 4.02
   + VLA-Cache 90.0% 90.0% 85.0% 73.3% 84.6% 1.303 51.85 4.21